# Enhancing Edge Detection by Texture Handling Architecture and Noiseless Training Data

### v4

@@ 4 Experiment
This section is dedicated to experiments. We begin by outlining the dataset settings and training methods, followed by a discussion of the evaluation metrics. Finally, we provide the experimental results and ablation studies.
@@ 4.1.1 Dataset
Our experiments were conducted on four datasets: BRIND[22], UDED[24], MDBD[26], and BIPED[25]. Since BRIND provides multiple annotations for each image, we consolidate all labels into a single binary edge map [^2] We merge all available ground-truths into a unified binary map, marking a pixel as an edge if it is labeled as such in any of the annotations.. For MDBD, we use annotation 3 as the ground-truth [^3] Annotation 3 is chosen arbitrarily; alternative combinations of the six available annotations were tested, but the ground-truth quality was found to be unsatisfactory.. UDED and BIPED each provide a single label per image, which is used as the ground-truth. Notably, among these datasets, UDED exhibits the highest annotation quality, while BRIND offers the most diverse image collection.
@@ 4.1.2 Data Augmentation
To enhance data diversity, training images are recursively split in half until their height and width are both below 640 pixels. Each resulting sub-image undergoes eight augmentation transformations: rotations of 0o, 90o, 180o, and 270o, combined with horizontal flips. Furthermore, to improve performance on noiseless data, the original labels themselves are added to the training set, with their ground-truths set to their own values.
@@ 4.1.3 Training Detail
All models are trained using the Adam optimizer with a weight decay of $10^{-8}$. The initial learning rate is set to $10^{-4}$ and is reduced by a factor of 10 every 50 epochs. The batch size is fixed at 8, and the balance parameter in the loss function is set to $\lambda=1.1$ (please refer to Equation (4)). During training, images are randomly cropped to 320$\times$320 pixels, and the dataset is refreshed every five epochs. To avoid resizing operations during training, the two UDED images with widths below 320 pixels are designated as test images. The number of training epochs is set as follows: BRIND: 100 epochs, UDED: 150 epochs, MDBD: 100 epochs, BIPED: 75 epochs.
@@ 4.2 Benchmark
4.2.1 Evaluation benchmark
To evaluate model performance, we adopt the widely used methodology from [7], which computes ODS, OIS, and AP under an error tolerance distance. However, previous studies have employed different error tolerance distances: approximately 4.3 pixels for BRIND [28], 11.1 pixels for BIPED [19], and up to 16.3 pixels for MDBD [26]. Many researchers retain the default evaluation setting, where the error tolerance distance is a fixed ratio of the image’s diagonal length, inadvertently leading to inconsistent benchmarks [^4] For instance, in BRIND (image resolution: $321\times 481$), an error tolerance ratio of 0.0075 corresponds to approximately 4.3 pixels, whereas in BIPED (resolution: $720\times 1280$), the same ratio allows for 11.1 pixels..
Using a diagonal-length ratio as the error tolerance criterion is problematic because it relaxes the evaluation standard as image resolution increases. For example, in the BSDS500 dataset (resolution: 321$\times$481), a ratio of 0.0075 results in a 4.3-pixel tolerance, whereas concatenating four such images into a single 642$\times$962 image effectively doubles the tolerance to 8.6 pixels. Consequently, artificially enlarging images can inflate evaluation scores without improving actual ED performance. To ensure a resolution-independent evaluation, we propose setting the error tolerance distance as a fixed pixel value rather than a proportional ratio.
Moreover, the choice of an appropriate error tolerance threshold for edge detection tasks has been largely overlooked. We argue that models should be evaluated using a low error tolerance distance, since poor performance under a strict threshold but high scores under a relaxed one indicate that erroneous predictions contribute disproportionately to the final score [^5] For example, a prediction missing a single edge pixel should score higher than one missing the same pixel while introducing an additional false edge elsewhere. However, a loose error tolerance threshold can increase the likelihood of the latter receiving a higher score, as the erroneous edge pixel may still fall within the tolerance range..
However, since edge annotations in current datasets are manually labeled, they often exhibit a one-pixel deviation due to human annotation inconsistencies. As illustrated in Fig. 4, edges are typically marked as one pixel wide, even though true edges should be two pixels wide. This means that ground-truth labels may vary by up to one pixel horizontally, vertically, or diagonally. Therefore, while a zero-pixel error tolerance represents the strictest possible standard, a one-pixel tolerance is more realistic given the labeling inconsistencies, whereas allowing more than one pixel of tolerance risks overestimating model accuracy. Based on these considerations, we propose using a one-pixel threshold for rigorous evaluation.
@@ 4.2.2 Evaluation Settings
All models are evaluated on full-resolution test images [^6] Unless explicitly stated, noiseless data are excluded from testing.. The standard evaluation algorithm from [7] is used, with the strictest error tolerance setting: one pixel in the diagonal direction (approximately 1.42–1.99 pixels in horizontal and vertical directions). The corresponding error tolerance distances are set as follows: BRIND: 0.003, MDBD and BIPED: 0.001, UDED (which contains images of varying resolutions): 0.004, ensuring a minimum of one-pixel tolerance for the smallest image ($250\times 361$ resolution).
As noted in [39], many models achieve SOTA performance on a specific dataset at the cost of generalization, leading to significant performance degradation when tested on different datasets [^7] This issue can arise even when only the training-test partitioning of the same dataset is altered.. To ensure a comprehensive and reliable evaluation, we conduct experiments on multiple datasets with various dataset partitions. We use the notation ’Dataset-$P_{k}$-a-b-Ec’ to index dataset partitions, where: Dataset refers to the dataset name, $P_{k}$ denotes the $k$-th partition, $a$ and $b$ represent the number of training and test images, respectively, and $c$ indicates the number of training epochs. For example, ’BRIND-$P_{1}$-250-250-E100’ refers to experiments conducted on the BRIND dataset using the first partition, where 250 images are used for training, 250 for testing, and the model is trained for 100 epochs.
@@ 4.2.3 Comparison Settings
The SDPED model is compared against several methods: HED[14], RCF[21], RIND [22], BDCN[23], and Dexi[19]. Since the SDPED architecture can be configured with different numbers of CSDB blocks, we denote a model with $n$ CSDB blocks as SDPED_$\_n$, with SDPED_$\_7$ being the standard for comparison. Moreover, for fairness, all models are trained under identical conditions, including the same dataset, optimizer, normalization strategy, and prediction-saving procedure.
Table 1:   Results on BRIND with error toleration 0.003 (1 pixel on the diagonal): The best results are highlighted in blue, while the second-best results are marked in red. In all three experiments and across all three benchmarks, our model (SDPED_$\_7$) achieves the best results. The improvements compared to the second-best results (for each partition and benchmark, regardless of the previous model) are also listed. On average, our model improves scores by approximately $2.7\%$ in ODS, $3.1\%$ in OIS, and $6.9\%$ in AP.
@@
HED (2015) 0.639 0.651 0.575 0.657 0.669 0.597 0.663 0.675 0.605
RCF (2017) 0.643 0.653 0.573 0.663 0.672 0.598 0.665 0.674 0.603
RIND (2021) 0.624 0.634 0.509 0.651 0.661 0.610 0.649 0.662 0.590
BDCN (2022) 0.639 0.652 0.567 0.645 0.659 0.576 0.662 0.673 0.587
Dexi (2023) 0.639 0.648 0.548 0.652 0.663 0.572 0.663 0.674 0.585
SDPED_$\_7$(Ours) 0.660 0.672 0.613 0.677 0.693 0.643 0.687 0.698 0.658
@@
SDPED_$\_5$(Ours) 0.660 0.672 0.616 0.676 0.692 0.645 0.683 0.693 0.649
SDPED_$\_3$(Ours) 0.654 0.668 0.617 0.674 0.687 0.647 0.684 0.695 0.658
Table 2:   Results on UDED with an error tolerance of 0.004 (equivalent to 1 pixel along the diagonal for the lowest-resolution image): The table is structured as Table 1. Across all three experiments and benchmarks, our model consistently achieves the best performance. The improvements over the second-best results are also reported. On average, our model (SDPED_$\_7$) improves scores by approximately $5.2\%$ in ODS, $4.4\%$ in OIS, and $2.4\%$ in AP.
@@
HED (2015) 0.714 0.732 0.687 0.709 0.733 0.673 0.639 0.675 0.590
RCF (2017) 0.737 0.775 0.713 0.746 0.763 0.724 0.637 0.703 0.609
RIND (2021) 0.416 0.425 0.071 0.375 0.381 0.046 0.716 0.747 0.678
BDCN (2022) 0.767 0.787 0.763 0.772 0.803 0.757 0.689 0.762 0.686
Dexi (2023) 0.771 0.784 0.766 0.767 0.793 0.779 0.701 0.750 0.731
SDPED_$\_7$(Ours) 0.812 0.828 0.792 0.807 0.830 0.791 0.757 0.798 0.748
@@
SDPED_$\_5$(Ours) 0.811 0.824 0.795 0.806 0.827 0.796 0.756 0.793 0.744
SDPED_$\_3$(Ours) 0.807 0.822 0.786 0.803 0.825 0.798 0.753 0.791 0.751
Figure 5:  Sample predictions on BRIND and UDED, all from partition $P_{1}$ and displayed without NMS. The results illustrate that our model generates more visually coherent and reliable predictions.
4.3 Experiment Results
The evaluation results on BRIND, UDED datasets are presented in Table 1 and 2, while results on MDBD and BIPED2 are exhibited in Table 5 and 6, in supplemental materials. Additionally, some predicted images are displayed in Fig 5. Our models achieve SOTA performance compared to previous approaches.
On BRIND, our model improves the score by approximately $2.7\%$ in ODS, $3.1\%$ in OIS, and $6.9\%$ in AP on average. On UDED, the improvements are approximately $5.2\%$ in ODS, $4.4\%$ in OIS, and $2.4\%$ in AP [^8] Notably, RIND exhibits inconsistent performance across different test cases. While some results ($P_{1}$ and $P_{2}$) are significantly subpar, the other ($P_{3}$) performs better. This instability is evident in our retraining experiments: when retrained three times on $P_{2}$, RIND consistently yielded low AP scores. Similarly, retraining it on $P_{1}$ three times resulted in two low-AP outcomes. In contrast, RIND was retrained on $P_{3}$ only once, where it performed comparatively better. These results suggest that RIND may lack stability across different scenarios.. On MDBD, considering error tolerances of 0.001 (1.5 pixels) and 0.003 (4.4 pixels), our model improves performance by about $4.8\%/2.5\%$ in ODS, $4.8\%/2.9\%$ in OIS, and $22.5\%/14.7\%$ in AP, respectively. On BIPED2, our model significantly outperforms previous ones in terms of AP (improving by approximately $11.8\%$ on average), and meanwhile achieves comparable results in ODS and OIS, with differences of less than $0.3\%$ on average—regardless of whether our model or previous ones obtained the best scores.
@@ 4.4 Ablation studies
We conduct ablation studies on BRIND and UDED to evaluate the effectiveness of the skipping dense structure and the multiple-layer fusion block. Table 3 demonstrates that both components contribute positively to the performance.
Table 3: Ablation study of SDPED_$\_7$ on BRIND and UDED: No skipping refers to removing all skipping connections in SDB (excluding residual ones), while Single fuse replaces the multiple-layer fusion block with a single $1\times 1$ convolution layer. No skipping with single fuse removes all skipping connections and applies a single $1\times 1$ convolution layer for fusion, both as above. The best results are highlighted in bold, demonstrating the effectiveness of the proposed structures.
@@ 5 Discussion
In this section, we discuss the effectiveness of the SDPED architecture, along with its limitations and potential directions for future work.
5.1 Reasons for the Effectiveness of SDPED
The performance improvements of SDPED models, especially in AP, can be attributed to several factors:
5.1.1 Avoiding Information Loss
Most previous models incorporate down-sampling techniques, such as max-pooling, to extract multi-scale features. However, these operations inevitably lead to information loss. For instance, a $2\times 2$ max-pooling operation merges information from four pixels, making it impossible to recover fine details, thereby increasing uncertainty and reducing precision. In contrast, SDPED models eliminate down-sampling altogether, preserving all original information and mitigating uncertainty.
Down-sampling has traditionally been employed in ED models based on the assumption that multi-scale information is crucial, as demonstrated in classical approaches [9]. However, in deep CNN architectures, the receptive field is already sufficiently large to capture multi-scale features, rendering explicit down-sampling redundant. Since down-sampling not only removes information but also increases uncertainty, eliminating it in SDPED enhances precision and improves overall model performance.
5.1.2 Enhancing Texture Differentiation
The performance of ED models largely depends on their ability to distinguish between textures and true edges. However, determining the most effective architectural design for this task remains challenging due to the inherent ambiguity in defining textures. Notably, the backbone of SDPED has demonstrated strong texture-handling capabilities in SR tasks. Given that both SR and ED require precise identification of texture regions, SDPED’s architecture is naturally well-suited for ED. [^9] In SR tasks, the objective is to identify texture regions and enhance them, whereas, in ED tasks, the goal is to detect texture regions and suppress them. A model that excels in SR is likely to be effective in ED, as both tasks rely on accurate texture localization.
5.1.3 Preserving the Strongest Responses

### v5

@@ 4 Experiment
This section presents the experimental validation of our framework. We first detail the dataset configurations and training protocols, followed by a description of the evaluation benchmarks. Finally, we provide experimental results and ablation studies.
@@ 4.1.1 Dataset
Our experiments were conducted on four datasets: BRIND[22], UDED[24], MDBD[26], and BIPED[25]. BRIND provides multiple annotations per image, from which we synthesize a unified binary ground-truth by marking a pixel as an edge if it is identified by any annotator. For MDBD, we use annotation 3 as the ground-truth [^2] Annotation 3 is chosen arbitrarily; alternative combinations of the six available annotations were tested, but the resulting ground-truths were found to be less reliable.. UDED and BIPED each provide a single label per image, which is used as the ground-truth. Notably, among these datasets, UDED exhibits the highest annotation quality, while BRIND offers the most diverse image collection.
@@ 4.1.2 Data Augmentation
To enhance data diversity, training images are recursively split in half until their height and width are both below 640 pixels. Each resulting sub-image undergoes eight augmentation transformations: rotations of 0o, 90o, 180o, and 270o, combined with horizontal flips. Furthermore, to implement the ideal prior guidance discussed in Section 3.2, we augment the training set by incorporating ground-truth edge maps as noiseless input samples, where each label serves as both the input and the target.
@@ 4.1.3 Training Detail
All models are trained using the Adam optimizer with a weight decay of $10^{-8}$. The initial learning rate is set to $10^{-4}$ and is reduced by a factor of 10 every 50 epochs. The batch size is fixed at 8, and the balance parameter in the loss function is set to $\lambda=1.1$ (please refer to Equation (4)). During training, images are randomly cropped to $\times320$ pixels, and the dataset is refreshed every five epochs. To avoid resizing operations during training, the two UDED images with widths below 320 pixels are designated as test images. The number of training epochs is set as follows: BRIND: 100 epochs, UDED: 150 epochs, MDBD: 100 epochs, BIPED: 75 epochs.
@@ 4.2 Benchmark
Evaluation benchmarks are clarified in this subsection.
@@ 4.2.1 Evaluation Settings
All models are evaluated on full-resolution test images [^3] Unless explicitly stated, noiseless data are excluded from testing.. The standard evaluation algorithm from [7] is used, under the strictest 1-pixel criterion (approximately 1.42–1.99 pixels in horizontal and vertical directions) proposed in Section 3.3. The maxDist parameter is calibrated for each dataset (BRIND: 0.003, MDBD/BIPED: 0.001, UDED: 0.004 (1-pixel tolerance for the smallest image)) to ensure a consistent 1-pixel error tolerance in the diagonal direction across varying resolutions.
As noted in [42], many models achieve SOTA performance on a specific dataset at the cost of generalization, leading to significant performance degradation when tested on different datasets [^4] This issue can arise even when only the training-test partitioning of the same dataset is altered.. To ensure robustness and reduce partition-specific bias, we conduct experiments on multiple datasets with various dataset partitions. We use the notation ’Dataset-$P_{k}$-a-b-Ec’ to index dataset partitions, where: Dataset refers to the dataset name, $P_{k}$ denotes the $k$-th partition, $a$ and $b$ represent the number of training and test images, respectively, and $c$ indicates the number of training epochs. For example, ’BRIND-$P_{1}$-250-250-E100’ refers to experiments conducted on the BRIND dataset using the first partition, where 250 images are used for training, 250 for testing, and the model is trained for 100 epochs.
@@ 4.2.2 Comparison Settings
The SDPED model is compared against several methods, including CNN-based and transformer-based: HED[14], RCF[21], RIND [22], BDCN[23], Dexi[19], and EdgeNat[36]. Since the SDPED architecture can be configured with different numbers of CSDB blocks, we denote a model with $n$ CSDB blocks as SDPED_n\_n, with SDPED_7\_7 being the standard for comparison. To ensure a rigorous and fair comparison, all models are retrained from scratch under identical conditions, utilizing the same unified training-test partitions, optimizer settings, and evaluation pipeline [^5] Models are trained with Adam with 0.0001 learning rate and decay by multiplying $\frac{1}{10}$ after 50 epochs. The number of batches is set to 8..
4.3 Experiment Results
The quantitative results on the BRIND, UDED, MDBD, and BIPED2 datasets are summarized in Tables 1 to 5, with qualitative visual comparisons provided in Figure 5. Following the evaluation protocol, the best and second-best performances are highlighted in blue and red, respectively. The bottom row of each table quantifies the performance margin achieved by SDPED7 relative to the previous SOTA.
Our framework obtains consistent improvements compared to previous methods, with particularly notable gains in AP, which is more sensitive to localization fidelity. On BRIND, SDPED7 achieves average gains of approximately 2.7% in ODS, 3.1% in OIS, and 6.9% in AP. On UDED, the improvements are approximately 5.2% in ODS, 4.4% in OIS, and 2.4% in AP [^6] It is worth noting that certain baseline models may exhibit significant performance fluctuations across different dataset partitions; while RIND performing reasonably on $P_{3}$, it yielded significantly lower scores on $P_{1}$ and $P_{2}$. When retraining RIND three times on $P_{2}$, it consistently yielded low scores, while retraining it on $P_{1}$ three times resulted in twice unsatisfactory outcomes. This empirical observation indicates that certain architectures may be sensitive to data partitioning, further motivating the use of multi-partition evaluations. This issue also appears on the dataset MBDB.. On MDBD, under both the strictest (0.001 or 1-pixel) and relaxed (0.003 or 4.4 pixels) error tolerances, our model demonstrates superior improvements in AP of about 22.5% and 14.7%, respectively, while also improving ODS/OIS by about $4.8\%/4.8\%$ and $2.5\%/2.9\%$ on the two criteria, respectively. On BIPED2, while SDPED7 achieves comparable ODS and OIS results (within 0.3% margin on average), it delivers a significant 11.8% improvement in AP.
Table 1:    Results on BRIND with error toleration 0.003 (1 pixel on the diagonal): Across all the experiments and benchmarks, our model (SDPED_7\_7) achieves the best results. On average, our model improves scores by approximately $2.7\%$ in ODS, $3.1\%$ in OIS, and $6.9\%$ in AP.
@@ 4.3 Experiment Results
HED 0.639 0.651 0.575 0.657 0.669 0.597 0.663 0.675 0.605
RCF 0.643 0.653 0.573 0.663 0.672 0.598 0.665 0.674 0.603
RIND 0.624 0.634 0.509 0.651 0.661 0.610 0.649 0.662 0.590
BDCN 0.639 0.652 0.567 0.645 0.659 0.576 0.662 0.673 0.587
Dexi 0.639 0.648 0.548 0.652 0.663 0.572 0.663 0.674 0.585
EdgeNat 0.570 0.577 0.379 0.577 0.584 0.385 0.590 0.597 0.396
SDPED_7\_7(Ours) 0.660 0.672 0.613 0.677 0.693 0.643 0.687 0.698 0.658
@@
SDPED_5\_5(Ours) 0.660 0.672 0.616 0.676 0.692 0.645 0.683 0.693 0.649
SDPED_3\_3(Ours) 0.654 0.668 0.617 0.674 0.687 0.647 0.684 0.695 0.658
Table 2:    Results on UDED with an error tolerance of 0.004 (1 pixel along the diagonal for the lowest-resolution image): Across all three experiments and benchmarks, our model consistently achieves the best performance. On average, our model improves scores by approximately $5.2\%$ in ODS, $4.4\%$ in OIS, and $2.4\%$ in AP.
@@
HED 0.714 0.732 0.687 0.709 0.733 0.673 0.639 0.675 0.590
RCF 0.737 0.775 0.713 0.746 0.763 0.724 0.637 0.703 0.609
RIND 0.416 0.425 0.071 0.375 0.381 0.046 0.716 0.747 0.678
BDCN 0.767 0.787 0.763 0.772 0.803 0.757 0.689 0.762 0.686
Dexi 0.771 0.784 0.766 0.767 0.793 0.779 0.701 0.750 0.731
EdgeNat 0.608 0.620 0.424 0.620 0.637 0.434 0.625 0.633 0.426
SDPED_7\_7(Ours) 0.812 0.828 0.792 0.807 0.830 0.791 0.757 0.798 0.748
@@
SDPED_5\_5(Ours) 0.811 0.824 0.795 0.806 0.827 0.796 0.756 0.793 0.744
SDPED_3\_3(Ours) 0.807 0.822 0.786 0.803 0.825 0.798 0.753 0.791 0.751
[…]
Partitions MDBD-$P_{1}$-60-40-E100 MDBD-$P_{2}$-80-20-E100 MDBD-$P_{3}$-80-20-E100
Benchmarks ODS OIS AP ODS OIS AP ODS OIS AP
HED 0.302 0.304 0.122 0.289 0.291 0.114 0.261 0.264 0.094
RCF 0.276 0.277 0.096 0.264 0.266 0.089 0.245 0.246 0.079
RIND 0.243 0.244 0.035 0.287 0.290 0.118 0.285 0.292 0.125
BDCN 0.277 0.278 0.104 0.270 0.273 0.097 0.255 0.257 0.088
Dexi 0.305 0.307 0.126 0.294 0.297 0.116 0.271 0.274 0.102
EdgeNat 0.251 0.255 0.075 0.218 0.224 0.065 0.216 0.220 0.064
SDPED_7\_7 0.327 0.330 0.165 0.309 0.312 0.148 0.291 0.297 0.139
[…]
Partitions MDBD-$P_{1}$-60-40-E100 MDBD-$P_{2}$-80-20-E100 MDBD-$P_{3}$-80-20-E100
Benchmarks ODS OIS AP ODS OIS AP ODS OIS AP
HED 0.565 0.570 0.425 0.550 0.554 0.396 0.515 0.520 0.356
RCF 0.538 0.540 0.356 0.525 0.527 0.339 0.492 0.495 0.307
RIND 0.455 0.460 0.091 0.555 0.564 0.438 0.531 0.550 0.445
BDCN 0.542 0.545 0.377 0.533 0.536 0.362 0.514 0.515 0.339
Dexi 0.571 0.576 0.433 0.558 0.563 0.407 0.521 0.529 0.374
EdgeNat 0.488 0.495 0.286 0.445 0.453 0.246 0.439 0.447 0.261
SDPED_7\_7 0.590 0.599 0.531 0.569 0.579 0.501 0.543 0.561 0.477
[…]
Partitions BIPED2-$P_{1}$-125-125-E75 BIPED2-$P_{2}$-175-75-E75
Benchmarks ODS OIS AP ODS OIS AP
HED 0.640 0.643 0.493 0.645 0.648 0.496
RCF 0.641 0.644 0.509 0.644 0.647 0.510
RIND 0.673 0.674 0.479 0.688 0.691 0.468
BDCN 0.654 0.657 0.519 0.667 0.670 0.530
Dexi 0.674 0.675 0.530 0.683 0.685 0.544
EdgeNat 0.574 0.576 0.362 0.587 0.589 0.380
SDPED_7\_7 0.675 0.678 0.599 0.684 0.687 0.608
[…]
BIPED2-$P_{3}$-200-50-E75 Average
HED 0.640 0.644 0.494 0.642 0.645 0.494
RCF 0.640 0.644 0.497 0.642 0.645 0.505
RIND 0.677 0.681 0.341 0.679 0.682 0.429
BDCN 0.655 0.659 0.512 0.659 0.662 0.520
Dexi 0.677 0.679 0.530 0.678 0.680 0.535
EdgeNat 0.559 0.562 0.343 0.573 0.576 0.362
SDPED_7\_7 0.673 0.680 0.588 0.677 0.682 0.598
[…]
Figure 5:    Sample predictions on BRIND and UDED, all from partition $P_{1}$ and displayed without NMS. The results illustrate that our model generates more visually coherent and spatially consistent predictions.
@@ 4.4 Ablation studies
We conducted ablation studies on the BRIND, UDED, and BIPED2 datasets to verify the contributions of our key architectural components: the CSDB structure and the multi-layer fusing block. As shown in Table 6, both components contribute positively and consistently to performance across datasets.
Table 6: Ablation study of SDPED_7\_7 on BRIND, UDED, and BIPED2: No skipping refers to removing all skipping connections in SDB (excluding residual ones), while Single fuse replaces the multiple-layer fusion block with a single $1\times 1$ convolution layer. No skipping with single fuse removes all skipping connections and applies a single $1\times 1$ convolution layer for fusion, both as above. The best results are highlighted in bold.
@@ 5 Discussion
In this section, we discuss plausible factors contributing to the effectiveness of the SDPED architecture, analyze its parameter efficiency and robustness across diverse scenarios, and identify limitations for future research.
5.1 Factors Contributing to the Effectiveness of SDPED
The observed performance improvements of SDPED, particularly in the AP metric, can be attributed to four core design philosophies:
5.1.1 Avoiding Information Loss by Removing Down-sampling
Unlike most previous ED models that rely on down-sampling (e.g., max-pooling) to extract multi-scale features, SDPED eliminates these operations entirely to maintain full-resolution processing throughout the pipeline. Down-sampling inherently introduces spatial uncertainty; for instance, a $2\times 2$ max-pooling operation collapses four pixels into one, which results in information loss and makes the precise recovery of fine edge locations highly challenging. While down-sampling was traditionally justified for expanding the receptive field [9], deep CNN architectures already possess a sufficiently large receptive field to capture multi-scale context. By removing down-sampling, SDPED preserves the original spatial information, thereby maximizing the localization precision required for high-frequency edge features.
5.1.2 Task-Adaptive Texture Differentiation
The primary challenge in ED is distinguishing between edges and textures. By performing a task-adaptive architectural transfer from image SR tasks, SDPED leverages the CSDB backbone modified from the DB, which has been extensively validated for its high-frequency feature handling capabilities. The logic behind this transfer is that both SR and ED share a fundamental sub-task: the precise localization of texture-rich regions. In SR, these regions are identified for enhancement, while in ED, they are identified for suppression. Our results suggest that an architecture capable of understanding texture for reconstruction can also be effectively adapted for texture suppression in detection tasks. Specifically, a model that excels in SR must accurately localize high-frequency textures to restore details; the same localization capability can be leveraged to differentiate these textures from edges.
5.1.3 Consistency-based Response Filtering